OpenStudy (anonymous):

If log (5x+20)=3, what is the value of x?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log(5x+20)=3 }\) \(\normalsize\color{blue}{ \log(5x+20)=3 \times \log 10 }\) \(\normalsize\color{blue}{ \log(5x+20)= \log 10^3 }\) \(\normalsize\color{blue}{ \log(5x+20)= \log 1000 }\)

OpenStudy (solomonzelman):

can you finish it ?

OpenStudy (anonymous):

no im kinda confused I thought when you had addition it turns into a product

OpenStudy (solomonzelman):

I am just giving you the fastest approach...

OpenStudy (anonymous):

so what do I do after?

OpenStudy (solomonzelman):

If \(\normalsize\color{blue}{ \log a= \log b }\) then \(\normalsize\color{blue}{ a= b }\)

OpenStudy (anonymous):

doesn't make any sense a)-19/5 b)-17/5 c)196 d)204

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

this what doesn't make sense, how can you ask questions about factoring (algebra 1) and about logarithms (trigonometry) at the same time ?

OpenStudy (anonymous):

ib mathematics

OpenStudy (anonymous):

x=196 (:

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

\[\log (5x+20)=3,\] means \[5x+20=10^3\]

OpenStudy (anonymous):

they say the same thing exactly that you can solve for \(x\) without much trouble

OpenStudy (solomonzelman):

I forgot to use just that relation -:(

OpenStudy (anonymous):

\[5x+20=1000\\ 5x=980\\ x=196\]